Suppose the banking system currently has $400 billion in reserves, the reserve requirement is 8 percent, and excess reserves amount to $5 billion. What is the level of deposits?

Selected Answer: d.

$4,937.5 billion

Answers: a.

$5,062.5 billion

  b.

$4,995 billion

  c.

$5,000 billion
Answer #1

Solution: $4,937.5 billion

Working:

Minimum Required Reserves = Current Reserves – Excess Reserves

Minimum Required Reserves = 400 – 5 = $395

R = (Min Required Reserves) / (Total Deposits)

Total Deposits = 395 / 0.08 = $4937.5